# BanStereoSet: A Dataset to Measure Stereotypical Social Biases in LLMs for Bangla

### v1

@@ Abstract
This study presents BanStereoSet, a dataset designed to evaluate stereotypical social biases in multilingual LLMs for the Bangla language. In an effort to extend the focus of bias research beyond English-centric datasets, we have localized the content from the StereoSet, IndiBias, and Kamruzzaman et al.’s (2024b) datasets, producing a resource tailored to capture biases prevalent within the Bangla-speaking community. Our BanStereoSet dataset consists of 1,194 sentences spanning 9 categories of bias: race, profession, gender, ageism, beauty, beauty in profession, region, caste, and religion. This dataset not only serves as a crucial tool for measuring bias in multilingual LLMs but also facilitates the exploration of stereotypical bias across different social categories, potentially guiding the development of more equitable language technologies in Bangladeshi contexts. Our analysis of several language models using this dataset indicates significant biases, reinforcing the necessity for culturally and linguistically adapted datasets to develop more equitable language technologies.
@@
Mahammed Kamruzzaman1, Abdullah Al Monsur1, Shrabon Das1, Enamul Hassan2, Gene Louis Kim1 1University of South Florida, 2North South University 1{kamruzzaman1, almonsur, das157, genekim}@usf.edu, 2enamul.hassan@northsouth.edu
@@ 1 Introduction
Bengali, with 233.7 million native speakers worldwide, ranks fifth worldwide. [^1] https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers Recent studies focusing on Bengali have primarily addressed gender and religious biases. For instance, Sadhu et al. (2024b) proposed a dataset concentrating on gender and religion biases through persona-based experiments, such as adopting the role of a typical Bengali person. Similarly, Das et al. (2023) evaluated these biases in Bengali. Sadhu et al. (2024a) found stereotypical emotional associations between genders in Bengali, using an emotion dataset.
@@ 2.1 Datasets and Translation
We translated the English sentences into Bangla using GPT-4, using few-shot learning to enhance the accuracy of the translations. After translating the sentences to Bangla, four native Bangla-speaking annotators, fluent in English as well, reviewed each translation for accuracy. Incorrect translations were appropriately modified. We included all intrasentence StereoSet examples and 250 samples from each bias type in GenAssocBias. We used only the caste-related examples from IndiBias.
@@ 2.5 Newly curated regional data
When curating regional data, we employed a human-LLM partnership. For both the attribute collection and sentence generation stages, we adhered to a criterion where data would only be retained if at least three out of the four annotators agreed on its appropriateness and contextual relevance.
We started with 17 major cities in Bangladesh and we used ChatGPT Brown et al. (2020) to generate sets of attributes for each city—10 stereotypical, 10 anti-stereotypical, and 10 unrelated. These attributes were intended to describe distinct regional characteristics. Subsequently, the annotators reviewed these attributes to ensure alignment with the actual characteristics of each region.
Following attribute validation, ChatGPT generated sentences that incorporated these attributes, leaving blanks specifically designed to reflect potential regional biases. This setup allowed the attributes to vary by region in the completed sentences. Once again, our team of annotators reviewed each sentence to confirm its contextual suitability and logical coherence. An example sentence of regional bias is presented in Figure 2(g).
The selected English sentences were translated into Bangla using GPT-4, and finally, we kept the Bangla-translated sentences adhering to the same translation validation process previously described in Section 2.1. City names and attributes used for region data curation are shown in Figure 4 in Appendix A. Each bias type in BanStereoSet contains the following number of examples—Race: 241, Gender: 178, Religion: 56, Profession: 206, Ageism: 134, Beauty: 130, Beauty Profession: 126, Caste: 60, and Region: 63.
@@ 3 Experimental Setup
We evaluated our Bangla dataset using four multilingual LLMs: GPT-4o, Mistral-7B, llama3.1-70B, and Gemma2-27B. We use three prompting templates: a baseline template and two paraphrases (see Table 2 for full templates). Our baseline prompting template was as follows: “Based solely on the context, select the appropriate word from the options provided, without considering the order in which the options come in, only provide the appropriate word. We do not require an explanation, just a single, plain Bangla text with the answer. All responses must be in Bangla.”. Following Gupta et al. (2023), we present our main results averaged across all three prompting templates.
To gauge how the results in Bangla compare to those in English (i.e., are the LLMs more biased, less biased, or similarly biased in Bangla compared to English?), we also ran the four models using the English version of the dataset. We use the modified English sentences (after removing all the pitfalls and all other adjustments for the Bangla context), corrected Bangla-translated sentences, and zero-shot learning. For further details on the models and results for each individual prompting template see Appendix B.
@@ 4 Results and Discussion
We present our main results for both Bangla and English in Table 1 in terms of stereotypical engagement/response rates, which indicate the percentage (%) of responses that align with stereotypical judgments, excluding unrelated engagements. [^3] We exclude unrelated engagements because in the StereoSet dataset unrelated terms are included solely to assess the overall quality of language models, not their biases. Conversely, in the GenAssocBias dataset, unrelated terms are used to examine the neutral engagement of LLMs, which is not influenced by bias.
@@ 5 Conclusion
We presented BanStereoSet, addressing a crucial gap in bias research by focusing on the Bangla language. This expansion broadens the scope beyond the commonly studied English-based datasets. Our findings show promise in LLM capability to handle bias in less common languages as well they do in English, but also show a stark gap between popular models in their ability to do so. This underscores the importance of culturally nuanced datasets for training equitable AI systems. We also found a key difference in the model’s ability to address caste and region bias between languages, emphasizing the necessity of extending bias research beyond English and encouraging the development of datasets that reflect diverse linguistic and cultural realities.
@@ 6 Limitations
The BanStereoSet dataset addresses significant gaps in bias evaluation for the Bangla language but presents several limitations that require consideration. Firstly, the dataset may not adequately represent regional variations in stereotypes throughout Bangladesh, as biases can vary markedly between different areas. Moreover, the dataset’s focus on binary gender (man and woman) representation restricts its ability to address biases concerning non-binary or gender-nonconforming identities. Challenges in translating stereotypes from English to Bangla may also lead to inaccuracies or cultural mismatches, although we try to address these translation issues in our annotation process, this is a good thing to keep in mind. Additionally, the relatively small size of certain categories, such as region bias, might limit the comprehensiveness of the bias evaluation. Although BanStereoSet encompasses a broad spectrum of biases, including race, gender, and religion, we didn’t include other crucial categories such as sexual orientation, socioeconomic status, or disability. Our dataset predominantly captures explicit biases, which may neglect more subtle or underlying biases. Furthermore, the effectiveness of language model evaluations using this dataset could be compromised by the models’ limited proficiency in Bangla, which might skew the results.
Acknowledgments
@@ Appendix B Experimental Setup and Prompting Details
We evaluate four major multilingual language models in this paper:
@@
4) Gemma2-27B, using the model from Ollama.
Additionally, we used tokenizers specific to each local LLM to properly format the prompts, as these models are instruction-tuned and require inputs to follow a particular structure. The tokenizers used are as follows:
@@
We also attempted to utilize Bangla-finetuned versions of Llama and Mistral, but the models’ responses were not reliable as they often produced results irrelevant to the topics and outside the given context. So, we excluded these Bangla-finetuned models.
To mitigate ordering bias, we followed the instructions described by Kamruzzaman and Kim (2024), instructing the models to ‘select the appropriate word from the options provided, without considering the order in which the options come in…’. We also set up our experiment in a way, where we randomly shuffle the order of choices presented to the LLMs.

### v2

@@ Abstract
This study presents BanStereoSet, a dataset designed to evaluate stereotypical social biases in multilingual LLMs for the Bangla language. In an effort to extend the focus of bias research beyond English-centric datasets, we have localized the content from the StereoSet, IndiBias, and Kamruzzaman et al.’s (2024b) datasets, producing a resource tailored to capture biases prevalent within the Bangla-speaking community. Our BanStereoSet dataset consists of 1,194 sentences spanning 9 categories of bias: race, profession, gender, ageism, beauty, beauty in profession, region, caste, and religion. This dataset not only serves as a crucial tool for measuring bias in multilingual LLMs but also facilitates the exploration of stereotypical bias across different social categories, potentially guiding the development of more equitable language technologies in Bangladeshi contexts. Our analysis of several language models using this dataset indicates significant biases, reinforcing the necessity for culturally and linguistically adapted datasets to develop more equitable language technologies. [^1] Our dataset is available at https://github.com/kamruzzaman15/BanStereoSet. 00footnotetext: This work has been accepted at ACL 2025.
@@
Mahammed Kamruzzaman1, Abdullah Al Monsur1thanks: *Equal contribution., Shrabon Das [^1] footnotemark: 1, Enamul Hassan2, Gene Louis Kim1 1University of South Florida  2North South University 1{kamruzzaman1, almonsur, das157, genekim}@usf.edu,  2enamul.hassan@northsouth.edu
@@ 1 Introduction
Bangla (also known as Bengali) is an Indic language belonging to the Indo-European language family. Bangla, with 237 million native speakers worldwide, ranks 7th worldwide [^2] Bangla is primarily spoken in Bangladesh, where it serves as the official language, and in the Indian state of West Bengal, as well as among diaspora communities worldwide. In this study, we focus solely on Bangladesh and its linguistic context, excluding other Bangla-speaking regions.. Recent studies focusing on Bengali have primarily addressed gender and religious biases. For instance, Sadhu et al. (2024b) proposed a dataset concentrating on gender and religion biases through persona-based experiments, such as adopting the role of a typical Bengali person. Similarly, Das et al. (2023) evaluated these biases in Bengali. Sadhu et al. (2024a) found stereotypical emotional associations between genders in Bengali, using an emotion dataset. More broadly, Mukherjee et al. (2023) analyzed socio-cultural prejudices across languages including Bangla by measuring the relative associations of word embeddings.
@@ 2.1 Datasets and Translation
We translated the English sentences into Bangla using GPT-4, incorporating few-shot examples to guide the translation process. After translating the sentences to Bangla, four native Bangla-speaking annotators, fluent in English as well, reviewed each translation for accuracy (see Appendix D for more details). Incorrect translations were appropriately modified. We included all intrasentence StereoSet examples and 250 samples from each bias type in GenAssocBias. We used only the caste-related examples from IndiBias.
@@ 2.5 Newly curated regional data
When curating regional data, we employed a human-LLM partnership. For both the attribute collection and sentence generation stages, we adhered to a criterion where data would only be retained if at least three out of the four annotators agreed on its appropriateness and contextual relevance. We started with 17 major cities in Bangladesh and we used GPT-4 to generate sets of attributes for each city—10 stereotypical, 10 anti-stereotypical, and 10 unrelated. These attributes were intended to describe distinct regional characteristics. Subsequently, the annotators reviewed these attributes to ensure alignment with the actual characteristics of each region. Following attribute validation, GPT-4 generated sentences that incorporated these attributes, leaving blanks specifically designed to reflect potential regional biases. This setup allowed the attributes to vary by region in the completed sentences. Once again, our team of annotators reviewed each sentence to confirm its contextual suitability and logical coherence. An example sentence of regional bias is presented in Figure 2(g). The selected English sentences were translated into Bangla using GPT-4, and finally, we kept the Bangla-translated sentences adhering to the same translation validation process previously described in Section 2.1. City names and attributes used for region data curation are shown in Figure 4 in Appendix A.
Bias Type Number of Examples
Race 241
Gender 178
Religion 56
Profession 206
Ageism 134
Beauty 130
Beauty Profession 126
Caste 60
Region 63
Table 1: Number of Examples per Bias Type in BanStereoSet.
The distribution of examples across different bias types within the BanStereoSet is detailed in Table 1. Each bias type in BanStereoSet contains the following number of examples—Race: 241, Gender: 178, Religion: 56, Profession: 206, Ageism: 134, Beauty: 130, Beauty Profession: 126, Caste: 60, and Region: 63.
@@ 3 Experimental Setup
We evaluated our Bangla dataset using four multilingual LLMs: GPT-4o, Mistral-7B, llama3.1-70B, and Gemma2-27B. We use three prompting templates: a baseline template and two paraphrases (see Table 3 for full templates). Our baseline prompting template was as follows: “Based solely on the context, select the appropriate word from the options provided, without considering the order in which the options come in, only provide the appropriate word. We do not require an explanation, just a single, plain Bangla text with the answer. All responses must be in Bangla.”. We present our main results averaged across all three prompting templates. To gauge how the results in Bangla compare to those in English (i.e., are the LLMs more biased, less biased, or similarly biased in Bangla compared to English?), we also ran the four models using the English version of the dataset. We use the modified English sentences (after removing all the pitfalls and all other adjustments for the Bangla context), corrected Bangla-translated sentences, and zero-shot learning. For further details on the models and results for each individual prompting template see Appendix B.
@@ 4 Results and Discussion
We present our main results for both Bangla and English in Table 2 in terms of stereotypical engagement/response rates, which indicate the percentage (%) of responses that align with stereotypical judgments, excluding unrelated engagements [^4] We exclude unrelated engagements because in the StereoSet dataset unrelated terms are included solely to assess the overall quality of language models, not their biases. Conversely, in the GenAssocBias dataset, unrelated terms are used to examine the neutral engagement of LLMs, which is not influenced by bias..
@@ 5 Conclusion
We presented BanStereoSet to address a critical gap in bias research by focusing on the Bangla language, expanding beyond the predominantly English-centric studies. Our findings highlight both the potential and limitations of LLMs in handling bias across languages, revealing significant disparities among models. This reinforces the ethical imperative of developing culturally informed datasets that ensure fairness and inclusivity in AI systems.
@@ 6 Limitations
Limitations of Non-Native Generated Text.
The BanStereoSet dataset addresses significant gaps in bias evaluation for the Bangla language but presents several limitations that require consideration. First, the reliance on non-native-speaker-generated text introduces potential discrepancies with real-world language usage. Despite efforts to mitigate this through human post-editing, the text may still lack the nuances and authenticity of native-speaker-generated data.
Dependence on English-Based Datasets.
Most of our work is based on previously created English datasets, which may not fully reflect contemporary language usage or the diverse contexts in which Bangla is employed on the internet but it allows direct comparison of LLM behaviors across languages.
Gender Representation Limitations.
Moreover, the dataset’s focus on binary gender (man and woman) representation restricts its ability to address biases concerning non-binary or gender-nonconforming identities.
Translating Stereotypes.
Challenges in translating stereotypes from English to Bangla may also lead to inaccuracies or cultural mismatches, although we try to address these translation issues in our annotation process, this is a good thing to keep in mind.
Category Imbalance and Bias Coverage.
Additionally, the relatively small size of certain categories, such as region bias, might limit the comprehensiveness of the bias evaluation. Although BanStereoSet encompasses a broad spectrum of biases, including race, gender, and religion, we didn’t include other crucial categories such as sexual orientation, socioeconomic status, or disability. Our dataset predominantly captures explicit biases, which may neglect more subtle or underlying biases.
Regional Limitations in Bangla Representation.
While Bangla is spoken beyond Bangladesh, including in West Bengal and global diaspora communities, our dataset focuses solely on the Bangladeshi context. This may overlook regional linguistic variations, cultural influences, and biases present in other Bangla-speaking populations. The Bengali language includes distinct dialects, with differences in pronunciation, vocabulary, and grammar between Bangladeshi and other Bangla-speaking people. Due to our limited knowledge of dialects spoken in other Bangla-speaking regions, we do not account for these variations in our bias evaluation.
LLM Proficiency in Bangla and Evaluation Challenges.
Furthermore, the effectiveness of language model evaluations using this dataset could be compromised by the models’ limited proficiency in Bangla, which might skew the results. Future research should consider incorporating data directly sourced from native speakers and real-world interactions to better align with actual language use and enhance the reliability of bias assessments in minority languages.
LLMs.
Our study is limited in its scope due to the restricted number of LLMs used for evaluation, which may not provide a comprehensive view of bias across different model architectures and training paradigms.
Acknowledgements
Mahammed Kamruzzman, Abdullah Al Monsur, Shrabon Das, and Gene Louis Kim were supported by the University of South Florida. We would like to thank Abdus Samad Tonmoy, University of Rajshahi, Bangladesh, for his initial annotation help. We also like to thank all the ARR anonymous reviewers for their valuable feedback on this paper.
@@ Appendix B Experimental Setup and Prompting Details
We evaluate four major multilingual language models in this paper. On model choosing, we try to choose both open-source and closed-source models in our experiments to balance our resource availability and cost:
@@
4) Gemma2-27B, using the model from Ollama. All models are used with their default hyperparameter settings. Additionally, we used tokenizers specific to each local LLM to properly format the prompts, as these models are instruction-tuned and require inputs to follow a particular structure. The tokenizers used are as follows:
@@
We also attempted to utilize Bangla-finetuned versions of Llama and Mistral, but the models’ responses were not reliable as they often produced results irrelevant to the topics and outside the given context. So, we excluded these Bangla-finetuned models. For few-shot data translation where we use GPT-4 we didn’t directly generate data using LLM, rather we just used GPT-4 to translate the data from English to Bangla only in the first place and then do the human annotation which helps to get data more efficiently.
To mitigate ordering bias, we followed the instructions described by Kamruzzaman and Kim (2024), instructing the models to ‘select the appropriate word from the options provided, without considering the order in which the options come in…’. We also set up our experiment in a way, where we randomly shuffle the order of choices presented to the LLMs. And then we run using three prompting templates as seen in Table 3.
@@ Appendix C Extended Results
Appendix D Data Annotation
In our study, we employed four native Bangla-speaking annotators from Bangladesh. Since our dataset is specifically designed for the Bangladeshi context (rather than other regions where Bangla is spoken, such as West Bengal in India), all annotators were selected from Bangladesh. Three of them are graduate students (PhD and Master’s level), while one is an undergraduate. All are fluent in English.
Most of these annotators specialize in ethics, bias, and fairness, giving them a strong understanding of biases. We conducted an extensive training session with them to discuss the annotation guidelines in detail. As part of their preparation, we instructed them to carefully read the paper titled ‘Stereotyping Norwegian Salmon: An Inventory of Pitfalls in Fairness Benchmark Datasets’ Blodgett et al. (2021). This helped them understand how pitfalls can arise during data creation and the possible ways to mitigate them.
The detailed annotation instructions for translating the data (except for our newly curated data, which follows a slightly different process) are provided in Figures 7, 8 and 9.
Figure 7: Instructions part 1.
Figure 8: Instructions part 2.
Figure 9: Instructions part 3.
Names.
We primarily focused on the Bangladeshi context, where the majority of the population identifies as Muslim. Therefore, the names included in our dataset predominantly reflect this demographic. However, we also included two names (Indrajit, Susmita) that are potentially Hindu to ensure some level of representational diversity within the dataset.